OpenStudy (unklerhaukus):

[Question from a DE textbook] \[\boxed{ \newcommand \dd [1] {\,\mathrm d#1} % infinitesimal \newcommand \de [2] {\frac{\mathrm d #1}{\mathrm d#2}} % first order derivative \newcommand \ps [1] {\left(#1\right)} % dynamic parentheses }\] Solve the following differential equation:\[x\de yx = y+\sqrt{x^2+y^2}\] Answer provided in the back of the book:\[y(x) = \frac12\ps{Ax^2-\frac1A}\] Answer from wolfram-alpha:\[y(x) = x\sinh\ps{\ln kx}\] Answer from my calculation (below):\[y(x) = kx^2-\sqrt{x^2+y^2}\] Why does my answer not match with either of the other answers? Have I made mistakes?

OpenStudy (unklerhaukus):

\begin{align} x\de yx &= y+\sqrt{x^2+y^2} \\ \de yx &= y/x +\sqrt{1+(y/x)^2} \\[-2ex] &&\text{let } y/x &= v \\ && y &= vx \\ && \de yx &= v+x\de vx \\[-2ex] v+x\de vx &= v +\sqrt{1+v^2} \\ x\de vx &= \sqrt{1+v^2} \\ \frac{\dd v}{\sqrt{1+v^2}} &= \frac{\dd x}x \\ \int\frac{\dd v}{\sqrt{1+v^2}} &= \int\frac{\dd x}x \\ \ln\ps{v+\sqrt{1+v^2}} &= \ln kx \\ v+\sqrt{1+v^2} &= kx \\ y/x+\sqrt{1+(y/x)^2} &= kx \\ y+\sqrt{x^2+y^2} &= kx^2 \\ y(x)&= kx^2-\sqrt{x^2+y^2} \end{align}

OpenStudy (unklerhaukus):

With a definition of the hyperbolic sine:\[\sinh \theta=\frac{e^\theta-e^{-\theta}}2\] The wolfram-alpha result: \[y(x)=x\sinh(\ln kx)\\ \qquad= x\ps{\frac{e^{\ln kx}-e^{-\ln kx}}2}\\ \qquad= x\ps{\frac{kx-\frac1{kx}}2}\\ \qquad= \frac12\ps{kx^2-\frac1k} \] can be show to equivalent the answer in the back of the book. But I don't know how to get my answer to be equilivalent

hartnn (hartnn):

your answer is equivalent to other answers, just isolate y :)

hartnn (hartnn):

(y-kx^2)^2 = x^2 + y^2 y^2 gets cancelled :)

OpenStudy (kainui):

This is such a dirty thing to put, a logarithm inside a hyperbolic trig function? \[\sinh (\ln a) = \frac{e^{\ln a}-e^{-\ln a}}{2} = \frac{1}{2}a + \frac{-1}{2} a^{-1}\] It's sorta awkward but maybe that's just a personal preference.

OpenStudy (unklerhaukus):

Got it now, thank you @hartnn ! \begin{align} y+\sqrt{x^2+y^2} &= kx^2 \\ y-kx^2 &= \sqrt{x^2+y^2} \\ y^2-2kyx^2+k^2x^4 &= x^2+y^2 \\ -2kyx^2+k^2x^4 &= x^2 \\ -2ky+k^2x^2 &= 1 \\ -2ky&= 1-k^2x^2 \\ y(x)&= \frac12\ps{kx^2-\frac1k} \end{align}

OpenStudy (unklerhaukus):

As far as I can see, the only advantage of the hyperbolic-logarithm form is that the constant only appears once, (rather than twice).